# SOMA: a novel sampler for exchangeable variables

### v1

@@ Abstract
The problem of sampling exchangeable random variables arises in many Bayesian inference tasks, especially in data imputation given a privatized summary statistics. These permutation-invariant joint distributions often have dependency structures that make sampling challenging. Component-wise sampling strategies, such as Metropolis-within-Gibbs, can mix slowly because they consider only comparing a proposed point with one component at a time. In this work, we propose a novel Single-Offer-Multiple-Attempts (SOMA) sampler that is tailored to sampling permutation invariant distributions. The core intuition of SOMA is that a proposed point unsuitable to replace one component might still be a good candidate to replace some other component in the joint distribution. SOMA first makes a singer offer, and then simultaneously considers attempts to replace each component of the current state with the single offer, before making the final acceptance or rejection decision. We provide an acceptance lower bound of SOMA and, using a coupling method, derive the convergence rate upper bound of SOMA in the two-dimensional case. We validate theoretical findings with numerical simulations, including a demonstration on differentially private Bayesian linear regression.
@@ 1 Introduction
Exchangability and permutation invariance arise in various problems in statistics and machine learning, such as de Finetti’s theorem (Diaconis and Freedman,, 1978), distributional symmetry (Austern and Orbanz,, 2022), graphons for exchangeable graphs (Caron and Fox,, 2017; Crane and Dempsey,, 2018), permutation-invariant neural networks (Lee et al.,, 2019; Bloem-Reddy et al.,, 2020; Kimura et al.,, 2024). The joint distribution of exchangeable random variables is invariant to permutations of the components. An important motivating example of our work is data imputation given noisy observations (Ju et al.,, 2022), especially when the noise process satisfies differential privacy (Dwork,, 2006). When missing or latent data are treated as random variables, they can often be viewed as exchangeable, and sampling from their conditional distributions allows us to complete the dataset for inference.
While exchangeability has deep theoretical foundations, for example, de Finetti’s theorem, Aldous–Hoover representations, etc., there’s relatively less work on how to construct or sample from general exchangeable distributions in practice, especially beyond simple or tractable cases.
To sample from distributions with intractable normalizing constants, we typically resort to Markov chain Monte Carlo (MCMC) methods. The Gibbs sampler (Gelfand and Smith,, 1990) is a popular MCMC algorithm to sample multivariate distributions. It updates the components in turn by sampling from its distribution conditional on the current values of the remaining components in the vectors. The Gibbs sampler is useful when the target distribution has tractable full conditional distributions when the joint distribution is not. When the full conditionals are intractable, one often uses the Metropolis-within-Gibbs sampler (MwG) (Gilks et al.,, 1995), which keeps the one-component-at-a-time structure of Gibbs but uses Metropolis-Hastings moves to approximate the conditional updates. A common Metropolis kernel for MwG has been the independent Metropolis-Hastings (IMH) (Tierney,, 1994; Liu,, 1996; Lee and Neal,, 2018) sampler. We refer to using IMH inside of MwG as independent-Metropolis-within-Gibbs (IMwG). While IMwG is suitable for sampling from multivariate distributions, it is not designed specifically to sample exchangeable variables.
To the best of our knowledge, there has not been a sampler exploits the symmetry of permutation invariant distributions in its design. In this work, we propose a novel Single-Offer-Multiple-Attempts (SOMA) sample leverages permutation invariance. Here is the core intuition of SOMA. During component-wise updates of IMwG, when a proposed point is unlikely to replace one component, it might still be likely to replace another component. Simultaneously considering multiple proposals might speed up the convergence of the sampler without adding significant computation cost. Unlike IMwG, SOMA transforms the IMH proposal into a state-dependent Metropolis kernel, and thus makes it more effective than conventional Gibbs style one-at-a-time updates. Our experiments indicate that as the number of components increases or as posterior correlation grows, SOMA remains highly efficient in avoiding stagnation, and results in better exploration of the state space than IMwG.
We also analyze the convergence rate of SOMA using the coupling method, and establish geometric ergodicity of SOMA for dimension 2. Theoretical analysis of the convergence rate of component-wise accept–reject-based Markov chains remains a challenging area (Brown and Jones,, 2022; Qin and Jones,, 2022; Qin et al.,, 2023), specially for dimensions higher than 2. Since our proof uses a coupling design, we believe that extending it to higher dimensions for a general permutation invariant distribution is challenging and would require stronger assumptions than what we currently use. We use numerical experiments to compare SOMA with IMwG up to 100 components (dimension $\approx 400$).
The remainder of this paper is structured as follows: In Section 2, we give examples of permutation invariant distributions and adapt MwG and IMH samplers to this setting. We present the SOMA algorithm in Section 3 and analyze its acceptance probability and convergence rate in Section 4. Experimental results are presented in Section 5.
@@ 2 Problem formulation and challenges
2.1 Permutation invariant distributions
Consider the problem of drawing samples from a multivariate target distribution on $\mathbb{X}^{n}$ with density $\pi(x_{1},\ldots,x_{n})$. In this work, we focus on densities $\pi$ exhibiting the following permutation invariance property.
Assumption 1 (Permutation invariance).
@@ 2.1 Permutation invariant distributions
For any permutation $\sigma$ of the group $\{1,2,\ldots,n\}$, we have $\pi(x_{\sigma(1)},x_{\sigma(2)},\ldots,x_{\sigma(n)})=\pi(x_{1},x_{2},\ldots,x_{n}).$ In other words, $(X_{1},\ldots,X_{n})\sim\pi$ is an exchangeable sequence of random variables.
Exchangability arises in various contexts of statistics. We give three illustrative examples here.
Example 1 (Bernoulli-Laplace diffusion model (Diaconis and Shahshahani,, 1987)).
Let $X_{i}\overset{iid}{\sim}\mathrm{Bernoulli}(p)$. Given the constraint $\sum_{i=1}^{n}X_{i}=s$, the conditional distribution of $X\mid\sum_{i=1}^{n}X_{i}=s$ is permutation invariant, and the density is
$\pi\left(x\mid\sum_{i=1}^{n}x_{i}=s\right)=\frac{1}{|\{x:\sum_{i=1}^{n}x_{i}=s\}|}\quad\text{for }\sum_{i=1}^{n}x_{i}=s.$ (1)
Example 2 (Data imputation given noisy summary statistic (Ju et al.,, 2022)).
Consider a confidential dataset $x=(x_{1},\ldots,x_{n})^{\top}$ consisting of $n$ data points. Suppose the data analysts can only observe a noisy summary statistic $s_{\mathrm{dp}}=s(x)+z$, where $s(x)$ is a summary statistic, and $z$ is the added noise. During data imputation, we can assume $x_{i}\overset{iid}{\sim}f(\cdot)$ and the observation model is $s_{\mathrm{dp}}\sim\eta(\cdot\mid s(x))$. The posterior density of $x$ given $s_{\mathrm{dp}}$ is thus given by
$\pi(x\mid s_{\mathrm{dp}})\propto\eta(s_{\mathrm{dp}}\mid x)\prod_{i=1}^{n}f(x_{i}).$ (2)
When the summary statistic $s(x)$ is invariant to the ordering of the components (e.g., the mean, median, or covariance), the data augmentation posterior $\pi(x\mid s_{\mathrm{dp}})$ is permutation invariant.
Example 3 (Bayesian PCA (Bishop,, 1998)).
Consider a dataset $v=(v_{1},\ldots,v_{m})^{\top}$ that each $v_{i}\in\mathbb{R}^{d}$ follows a conditional Gaussian distribution $p(v_{i}\mid u_{i})=\mathcal{N}(\Theta u_{i}+\mu,\sigma^{2}I_{d})$, where $I_{d}$ is the $d$-dimensional identity matrix, $\Theta=(\theta_{1},\ldots,\theta_{n})$ is a $d\times n$ matrix, and $u_{i}$ is a $n$-dimensional latent variable with prior $\mathcal{N}(0,I_{n})$. Let $C=\sigma^{2}I_{d}+\sum_{i=1}^{n}\theta_{i}\theta_{i}^{\top}$ and $S$ be the sample covariance matrix of $v$. Then the conditional density of $\Theta$ given the other parameters is $p(\Theta\mid\mu,\sigma^{2},v)\propto\exp[-m(\ln|C|+\mathrm{Tr}(C^{-1}S))/2]$, which is invariant to permutation of the columns $\theta_{i}$.
@@ 2.2 Gibbs, Metropolis, and Metropolis-within-Gibbs
When direct sampling from the joint distribution is difficult, a Markov chain Monte Carlo strategy is the Gibbs sampler (Geman and Geman,, 1984; Gelfand and Smith,, 1990). Gibbs sampling is ubiquitous in Bayesian modeling and it is particularly useful when one can sample from the full conditional distributions
@@
Here $x_{-i}$ includes all elements of $x$ except $x_{i}$, i.e., $x_{-i}=(x_{1},x_{2},\ldots,x_{i-1},x_{i+1},\ldots,x_{n}).$ Note that permutation invariance allows us to write the conditional as $\pi(x_{i}\mid x_{-i})$ instead of $\pi_{i}(x_{i}\mid x_{-i}).$ Gibbs samplers entails iterative updates according to $x_{i}\sim\pi(\cdot\mid x_{-i}).$ There are two main update rules for the Gibbs sampler (Andrieu,, 2016): systematic-scan (SysScan, also called deterministic-scan) and random-scan (RanScan). The former method sequentially updates all components once at a time with a fixed schedule $i=1,2,\ldots,n$, and the latter randomly selects a component $i$ to perform the conditional update in each iteration.
In many practical scenarios (e.g., Example 2), the full conditional distribution $\pi(x_{i}\mid x_{-i})$ is intractable and hence direct sampling is difficult. In that case, we must resort to approximated Gibbs samplers (Qin et al.,, 2023) such as the Metropolis-within-Gibbs sampler (Gilks et al.,, 1995), which uses a Metropolis-Hastings kernel to approximate each condition distribution.
A significant drawback of general-purpose, off-the-shelf methods is that they do not attempt to exploit the symmetry of permutation invariant distributions. One strategy to exploit permutation invariance is to first restrict sampling from the space of ordered sequences $\{x_{1}\leq x_{2}\leq x_{3}\leq\ldots\leq x_{n}\}$ and then obtain samples from the full space by permuting indices in the ordered sequences. As illustrated in Figure 1, it is unclear whether sampling from the constrained space is simpler. At the same time, accept-reject type of MCMC sampling from such a constrained space is challenging; see Lan and Kang, (2023) for a review on sampling from spaces with constraints like $c(x)=0$ and $c(x)\geq 0$.
Figure 1: Comparison of a permutation invariant density in the full space $\mathbb{R}^{2}$ (left) and in the constrained space $\{x\in\mathbb{R}^{2}:x_{1}\leq x_{2}\}$ (right).
Another implication of permutation invariance is that marginal distributions of $\pi$ are identical.
Lemma 1 (Marginal equivalence).
If the density $\pi(x_{1},\ldots,x_{n})$ is permutation invariant, then the marginal distributions $\pi(x_{i})$ are identical for all $i\in\{1,\ldots,n\}$.
The inherent symmetry from permutation invariance suggests a certain homogeneity among the coordinates. As a result, it is reasonable to use the same Metropolis kernel in each Gibbs step. In particular, we can use independent Metropolis-Hastings (Tierney,, 1994; Liu,, 1996; Lee and Neal,, 2018; Wang,, 2022), making the entire algorithm IMH within Gibbs (IMwG). In general, IMH uses proposal distributions that are independent of the current state. In the context of IMwG, when updating the $i$-th component, IMH proposes a new state $y\sim q(\cdot)$ from $\mathbb{X}$, independent of the current state $x_{i}$. Replacing $x_{i}$ by $y$ with probability
@@
leaves the (conditional) target distribution $\pi(x_{i}\mid x_{-i})$ invariant. If accepted, then the new state $x^{\prime}=[x_{-i},y]$ is the result of replacing the $i$-th component of $x$ by $y$ and keeping all other components. We will compare our proposed method with IMwG, under both RanScan and SysScan schedules. We refer to these two baselines as Ran-IMwG and Sys-IMwG, respectively.
3 SOMA: A Single-Offer-Multiple-Attempts sampler
We describe our novel Single-Offer-Multiple-Attempts (SOMA) sampler in this section. SOMA is motivated by the following observation about IMwG. When a proposed value $y$ to replace $x_{i}$ is rejected, it can still be a favorable candidate to swap out a different component $x_{j}$. In this sense, plain IMwG is not using the proposed samples efficiently.
Unlike the IMwG strategy of comparing $y\in\mathbb{X}$ with one component $x_{i}$ at a time, SOMA simultaneously considers a single proposed state $y$ as a candidate to replace multiple components. With a sample $y$, SOMA first randomly choose a component $1\leq i\leq n$ with probability proportional to
$w_{i}(y,x)=\frac{\pi([x_{-i},y])}{q(y)\prod_{j=1,j\neq i}^{n}q(x_{j})},\quad\text{for }i=1,2,\ldots,n.$ (5)
In other words, the probability of choosing the $i$-th component is $\mathbb{P}(I=i\mid y,x)=w_{i}(y,x)/\left(\sum_{i=1}^{n}w_{i}(y,x)\right)$, and we write the normalizing constant as $W(y,x)=\sum_{i=1}^{n}w_{i}(y,x).$ This completes the first stage of SOMA, which chooses an index $i$ based on the current state $x\in\mathbb{X}^{n}$ and a proposed component $y$. In the second stage, SOMA attempts to swap out $x_{i}$ and replace it by $y$ with probability

### v2

# SOMA: A Novel Sampler for Bayesian Inference from Privatized Data

@@ Abstract
Making valid statistical inferences from privatized data is a key challenge in modern analysis. In Bayesian settings, data augmentation MCMC (DAMCMC) methods impute unobserved confidential data given noisy privatized summaries, enabling principled uncertainty quantification. However, standard DAMCMC often suffers from slow mixing due to component-wise Metropolis-within-Gibbs updates. We propose the Single-Offer-Multiple-Attempts (SOMA) sampler. This novel algorithm improves acceptance rates by generating a single proposal and simultaneously evaluating its suitability to replace all components. By sharing proposals across components, SOMA rejects fewer proposal points. We prove lower bounds on SOMA’s acceptance probability and establish convergence rates in the two-component case. Experiments on synthetic and real census data with linear regression and other models confirm SOMA’s efficiency gains.
@@ 1 INTRODUCTION
Imputing data from noisy summary statistics is a foundational problem with broad applications, especially in privacy-preserving data analysis. In these applications, the data analysts cannot access the raw data directly, and only the noise-corrupted results from certain summary statistics queries are shared with them. By sampling from the conditional distribution of the missing raw data given observed noisy summary statistics, data imputation enables downstream tasks that require individual-level data. As the target distributions are multivariate and their densities can be evaluated only up to an unknown normalizing constant, data imputation typically requires Markov Chain Monte Carlo samplers.
Ju et al., (2022) proposes a data augmentation (Van Dyk and Meng,, 2001) strategy to make inference from privatized data. The data augmentation MCMC (DAMCMC) strategy has since also been adopted by (Guo et al.,, 2024; Awan et al.,, 2024) among others. DAMCMC consists of an inference step and an imputation step. This gives a chain that approximates the joint distribution of confidential data and parameters given observed noisy summary statistics. For the imputation step, Ju et al., (2022) propose using the independent Metropolis-Hastings (IMH) (Tierney,, 1994; Liu,, 1996; Lee and Neal,, 2018) sampler for each conditional update within a Gibbs sampler. We refer to this strategy as independent-Metropolis-within-Gibbs (IMwG). They show that the acceptance probability of IMwG proposed states are lower bounded, and connected to the magnitude of privacy noise. This makes IMwG suitable for the problem of interest. However, there remains an opportunity to design new sampling algorithms with a provably faster convergence than the IMwG method.
Our contributions are three folded.
  • First, we proposes a novel Single-Offer-Multiple-Attempts (SOMA) sampler that exploits a permutation-invariant property of the data imputation distribution. To the best of our knowledge, no existing method has exploited this property.
  • Second, we investigate the theoretical properties of SOMA and IMwG. We show that the acceptance probability of SOMA is larger than that of IMwG. For the case of $n=2$ components, we establish convergence rates of SOMA and IMwG by constructing a Markovian coupling. We also prove that SOMA becomes rejection-free as $n\to\infty.$
  • Finally, we apply SOMA to the problems of Bayesian linear regression given privatized data and Bayesian inference from privatized compositional data on American Time Use Survey (ATUS) data.
The remainder of this paper is structured as follows: Section 2 defines the data imputation problem, establishes the permutation-invariant property of the target distribution, and reviews the IMwG method and its limitations. We present the SOMA algorithm in Section 3 and analyze its acceptance probability and convergence rate in Section 4. Simulation and real data experiments are presented in Section 56. The appendix contains proofs and additional experiments.
@@ 2 PROBLEM FORMULATION AND CHALLENGES
The primary motivation of SOMA is to make inferences from observed data that have been corrupted to satisfy privacy guarantees.
2.1 Differential Privacy and Data Release
Consider a confidential dataset $X=(X_{1},\ldots,X_{n})^{\top}$ consisting of $n$ records, and each $X_{i}\in\mathbb{X}.$ When $X$ contains sensitive information, only sanitized results $S_{\mathrm{dp}}\in\mathbb{S}$ are released to the data users for analysis. Note that $\mathbb{S}$ need not have the same dimension with $\mathbb{X}$ or $\mathbb{X}^{n}$. When $S_{\mathrm{dp}}$ is a sanitized database, we have $S_{\mathrm{dp}}\in\mathbb{X}^{k}$, where $k$ is the number of sanitized samples. When $S_{\mathrm{dp}}$ is some privatized query result, $S_{\mathrm{dp}}$ has the same dimension as the query. The methodology we develop is applicable to both instances.
A data release mechanism $\eta(\cdot\mid X)$ is a conditional distribution of $S_{\mathrm{dp}}$ given $X$. This paper focuses on mechanisms that satisfy the definition of differential privacy (DP) from Dwork, (2006).
Definition 1 (Distance between datasets).
Let $x,x^{\prime}$ be two datasets of the same size. Define $d(x,x^{\prime})$ as the minimum number of components in which $x$ and $x^{\prime}$ differ, up to permutation of indices, i.e.,
$d(x,x^{\prime})=\min_{\sigma\in S_{n}}\|x_{\sigma}-x^{\prime}\|_{0},$
where we use $x_{\sigma}$ to represent the vector obtained from permuting indices from $x$. We call $x,x^{\prime}$ neighboring datasets if $d(x,x^{\prime})=1.$
Definition 2.
A mechanism $S_{\mathrm{dp}}\sim\eta(\cdot\mid X)$ satisfies $\epsilon$-DP, if, for all pairs of neighboring databases $(x,x^{\prime})$, we have
$\sup_{B\in\mathcal{B}}\frac{\eta(B\mid x)}{\eta(B\mid x^{\prime})}\leq\exp(\epsilon),$ (1)
where $\mathcal{B}$ is the measurable ball on $\mathbb{S}$ where the parameter $\epsilon>0$ is called the privacy loss budget. The ratio is defined to be 1 whenever the numerator and denominator are both 0.
The Laplace mechanism is a common technique for adding noise to achieve privacy. Given a query function $s:\mathbb{X}^{n}\to\mathbb{S}$, and query result $s(X)$, we have $S_{\mathrm{dp}}=s(X)+\nu$ where components in $\nu$ are independent and identical samples from a zero-centered Laplace distribution. The variance is determined by the sensitivity of $s$ as well as the privacy loss budget $\epsilon$. We give some examples of privatized query results here. These mechanisms will be covered by our experiments in Sections 5, 6, and B.1.
Example 1 (Perturbed histogram).
Let the confidential data domain be $\mathbb{X}=[0,1]^{r}$. We create $m$ bins in $\mathbb{X}$, and denote the $j$-th bin as $B_{j}$. Let $S_{j}$ be the number of $X_{i}$’s following in the $j$-th bin. Here $(S_{1},\ldots,S_{m})$ is a query result from $X$. Let $S_{\mathrm{dp},j}=S_{j}+\nu_{j}$ where $\nu_{j}$ are independent draws from $\mathrm{Laplace}(0,2/\epsilon).$ The perturbed histogram mechanism (Dwork,, 2006) is to release $S_{\mathrm{dp}}$. One can also generate synthetic data from perturbed histograms (Wasserman and Zhou,, 2010). Let $\tilde{S}_{j}=\max(0,S_{\mathrm{dp},j})$, $q_{j}=\tilde{S}_{j}/\left(\sum_{j=1}^{k}\tilde{S}_{j}\right)$. Sample $Z_{i}$ independently according to (unormalized) density $\sum_{j=1}^{m}q_{j}\mathbb{I}(z\in B_{j})$ for $i=1,\ldots,k$, gives a sanitized database $Z=(Z_{1},\ldots,Z_{k})$ of size $k$.
Example 2 (Data release for privatized linear regression).
Consider a regression setting with predictors $x_{i}\in\mathbb{R}^{p}$ and a response $y_{i}\in\mathbb{R}$. The confidential data is a concatenated matrix $Z$ with rows $(x_{i}^{\top},y_{i})$. To ensure finite sensitivity, each entry in $Z$ is first clamped to a pre-defined range. The summary statistics are the unique elements of the Gram matrix $Z^{\top}Z$. A privatized version is released by adding iid Laplace noise, scaled to the sensitivity of these statistics with each unique element (Bernstein and Sheldon,, 2019).
Example 3 (Private release of compositional data).
Let $\mathbb{X}=\{0\leq x_{j}\leq 1,\sum_{i=1}^{p}x_{j}=1\}$. The summary statistic is constructed from transformed data. First, each record $X_{i}$ is processed by element-wise clamping to an interval $[a,1]$, followed by applying the logarithm. The sum of these transformed vectors, $S_{\log}=\sum_{i=1}^{n}\log([X_{i}]_{a}^{1})$. A privatized version is then released as $S_{\mathrm{dp}}=S_{\log}+\nu$, where the components of the noise vector $\nu$ are iid draws from a $\mathrm{Laplace}(0,-p\log(a)/\epsilon)$ distribution (Guo et al.,, 2024).
2.2 Inference from Privatized Data and Data Augmentation
Our goal is to make inferences about the data-generating process $X_{i}\overset{iid}{\sim}f_{\theta}(\cdot)=f(\cdot\mid\theta)$ based on privatized data $S_{\mathrm{dp}}$. Sometimes, a data analyst can only access some query result $s(x)$ about the confidential dataset, and the data maintainer might only share a privatized version of it, resulting in $s_{\mathrm{dp}}=s(x)+z$, where $z$ is some additive noise independent of $s(x).$ The observation density $\eta(s_{\mathrm{dp}}\mid s(x))$ depends on $s_{\mathrm{dp}}-s(x)$ and known privacy parameters.
In this setting, we need to make inferences about $\theta$ based on $s_{\mathrm{dp}}$, without direct access to $x$. This article focuses on Bayesian inference, where the data analyst describes their prior belief about $\theta$ with some prior distribution $p(\theta).$ Ju et al., (2022) proposes a data augmentation strategy (Van Dyk and Meng,, 2001) using a Markov Chain Monte Carlo algorithm to approximate the joint posterior distribution
$\pi(\theta,x\mid s_{\mathrm{dp}})\propto p(\theta)\cdot\prod_{i=1}^{n}f(x_{i}\mid\theta)\cdot\eta(s_{\mathrm{dp}}\mid s(x)).$ (2)
From (2), the marginal distribution of $\theta$ is $\pi(\theta\mid s_{\mathrm{dp}})$. The algorithm alternates between two steps:
  • Inference step: (approximate) sampling from the conditional distribution $\pi(\theta\mid x,s_{\mathrm{dp}});$
  • Imputation step: (approximate) sampling from the conditional distribution
    $\pi(x\mid\theta,s_{\mathrm{dp}})\propto\eta(s_{\mathrm{dp}}\mid s(x))\prod_{i=1}^{n}f(x_{i}\mid\theta)$ (3)
When focusing on the imputation step, $\theta$ and $s_{\mathrm{dp}}$ are fixed, and we can write $\pi(x\mid\theta,s_{\mathrm{dp}})$ as $\pi(x)$ to simplify the notation. This work designs a sampler for the imputation step by exploiting the permutation invariant property of (3).
Definition 3 (Permutation invariance).
@@ 2.2 Inference from Privatized Data and Data Augmentation
A distribution $\pi(x)=\pi(x_{1},\ldots,x_{n})$ is permutation invariant if, for any permutation $\sigma$ of the group $\{1,2,\ldots,n\}$, we have $\pi(x_{\sigma(1)},x_{\sigma(2)},\ldots,x_{\sigma(n)})=\pi(x_{1},x_{2},\ldots,x_{n}).$ In other words, $(X_{1},\ldots,X_{n})\sim\pi$ is an exchangeable sequence of random variables.
Lemma 4.
Whenever the query $s(x)$ is permutation invariant, the conditional distribution (3) is also permutation invariant.
Proof.
Since $s(x)$ and $\prod_{i=1}^{n}f(x_{i}\mid\theta)$ are both permutation invariant, their product (3) is also permutation invariant.
Many queries like the mean, median, empirical distribution, or covariance are invariant to the ordering of the components. In particular, Examples 1, 3 and 2 all lead to permutation-invariant imputation steps. As a result, the imputation step is to sample from a permutation-invariant distribution. Previous work (Ju et al.,, 2022; Awan et al.,, 2024; Xiong et al.,, 2025) have not leveraged this property of $\pi(x\mid s_{\mathrm{dp}},\theta).$
The methodology we propose is applicable to any multivariate distribution that is permutation invariant, which arises in various other problems in statistics and machine learning.
@@ 2.3 Gibbs, Metropolis, and Metropolis-within-Gibbs
When direct sampling from the joint distribution $\pi(x)$ is difficult, a Markov chain Monte Carlo strategy is the Gibbs sampler (Geman and Geman,, 1984; Gelfand and Smith,, 1990). Gibbs sampling is ubiquitous in Bayesian modeling, and it is particularly useful when one can sample from the full conditional distributions
@@
Here $x_{-i}$ includes all components of $x$ except the $i$-th component, i.e., $x_{-i}=(x_{1},\ldots,x_{i-1},x_{i+1},\ldots,x_{n}).$ Note that permutation invariance allows us to write the conditional as $\pi(x_{i}\mid x_{-i})$ instead of $\pi_{i}(x_{i}\mid x_{-i}).$
Gibbs samplers entails iterative updates according to $x_{i}\sim\pi(\cdot\mid x_{-i}).$ There are two main update schedules for the Gibbs sampler (Andrieu,, 2016): systematic-scan (SysScan, also called deterministic-scan) and random-scan (RanScan). The former method sequentially updates all components once at a time with a fixed schedule $i=1,2,\ldots,n$, and the latter randomly selects a component $i$ to perform the conditional update in each iteration.
In many practical scenarios and especially for data imputation tasks the full conditional distribution $\pi(x_{i}\mid x_{-i})$ is intractable and hence direct sampling is difficult. In that case, we must resort to approximated Gibbs samplers (Qin et al.,, 2025) such as the Metropolis-within-Gibbs sampler (Gilks et al.,, 1995), which uses a Metropolis-Hastings kernel to approximate each conditional distribution.
The inherent symmetry from permutation invariance of (3) suggests a certain homogeneity among the full conditionals (4). Ju et al., (2022) notices that it is reasonable to use the same Metropolis kernel in each Gibbs step. In fact, they propose to use the independent Metropolis-Hastings (Tierney,, 1994; Liu,, 1996; Lee and Neal,, 2018; Wang,, 2022) sampler to approximate (4), making the entire algorithm IMH within Gibbs (IMwG). In general, IMH uses proposal distributions that are independent of the current state. In the context of IMwG, when updating the $i$-th component, IMH proposes a new state $y\sim q(\cdot)$ from $\mathbb{X}$, independent of the current state $x_{i}$. Replacing $x_{i}$ by $y$ with probability
@@
leaves the (conditional) target distribution $\pi(x_{i}\mid x_{-i})$ invariant. If accepted, then the new state $x^{\prime}=[x_{-i},y]$ is the result of replacing the $i$-th component of $x$ by $y$ and keeping all other components.
With full conditional (4) and proposing from the model $q(\cdot)=f(\cdot\mid\theta)$ simplifies (5) to
$\min\left(1,\frac{\eta(s_{\mathrm{dp}}\mid s([x_{-i},y]))}{\eta(s_{\mathrm{dp}}\mid s(x))}\right).$ (6)
This computation can be completed in $\mathcal{O}(1)$ time when the privacy mechanism satisfies a ‘recording additivity’ property (Ju et al.,, 2022, Assumption 2).
We will compare our proposed method with IMwG, under both RanScan and SysScan schedules. We refer to these two baselines as Ran-IMwG and Sys-IMwG, respectively.
3 A SINGLE-OFFER-MULTIPLE-ATTEMPTS SAMPLER
We describe our novel SOMA sampler in this section. SOMA is motivated by the following observation about IMwG. When a proposed value $y$ to replace $x_{i}$ is rejected, it can still be a favorable candidate to swap out a different component $x_{j}$. In this sense, plain IMwG is not using the proposed samples efficiently.
Unlike the IMwG strategy of comparing $y\in\mathbb{X}$ with one component $x_{i}$ at a time, SOMA simultaneously considers a single proposed state $y$ as a candidate to replace multiple components. With a sample $y$, SOMA first randomly chooses a component $I$ from $1,2\ldots,n$ to replace. The selection probability is proportional to weights $w_{i}$ given by
$w_{i}(y,x)=\eta(s_{\mathrm{dp}}\mid s([x_{-i},y])),\quad\text{for }i=1,\ldots,n.$ (7)
The probability of choosing the $i$-th component is then $\mathbb{P}(I=i\mid y,x)=w_{i}(y,x)/\left(\sum_{i=1}^{n}w_{i}(y,x)\right)$. Using weights (7) means that SOMA favors swapping out components that are less likely to have produced $s_{\mathrm{dp}}$. Notice that $w_{i}$ coincides with the denominator in (6).
In the second stage, given $I=i$, SOMA attempts to swap out $x_{i}$ and replace it by $y$ with probability